40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2384     Entry time: Wed Aug 7 17:53:46 2019
Author: anchal 
Type: DailyProgress 
Category: FSS 
Subject: Investigation of North TTFSS Box 

Measurement setup and analysis:

  • EOM and Laser PZT were disconnected and only the TTFSS box was being analyzed as a standalone circuit.
  • I used SR 785 to take transfer functions from 1 Hz to 102.4 kHz with 500 points, 1 settle cycle and 5 integration cycle. The template file used is in the code and data link below.
  • I used AG 4395A to take transfer functions from 103 kHz to 903 kHz with 1kHz IF bandwidth and 25 averages for some of the measurements. The template file is included.
  • 1st measurement was taken from Test1 input port at RF Board D0901894 to the OUT2 port on the common path in board D040105.
  • Remaining measurements are taken with Test1 input port terminated with 50 Ohm and input from Test2 port on the common path in board D040105 to various test points as marked on the plots.
  • I followed essentially same measurements on the zero model of the circuit from TTFSSzero package I recently created.
  • I then divided transfer functions to obtain testpoint to testpoint stagewise transfer functions, for both measured data and simulated data and compared the results.
  • In some cases where I felt data above 102.4 kHz is required, I used AG 4395A for that and concatenated the two results. Hence there is a step-change in measured transfer functions at 102.4 kHz in some cases.

Conclusions:

  • The first attached file, NorthFSSInvestigationMismatches.pdf has the plots which had a significant difference between measurements and simulations. Other stages are more or less working as expected.
  • Clearly something is off between test point 15 and test point 16. This is the third stage in PZT path where there is a notch and boost switch.
  • The boost is not working at all and our notch might be placed at a different place than simulations. However, simulations use 35 pF value for C46 on board D040105 which is just arbitrary and we don't really know the real tuned value.
  • From test point 16 to 18 too, there is significant difference after 10 kHz, but I'm not sure if this is an artifact of my measurement and simulation technique or not.
  • Finally, on the EOM Path, everything is good upto testpoint 12. It is right before the HV amplifier that there is a significant difference in what simulation suggests and what is actually happening.
  • I should maybe redo this last measurement with low excitation voltage to ensure nothing is railing.
  • Also, the appearance of a notch at 4 Hz is just very weird in this last simulation. It was not predicted by an earlier similar analysis using pyliso.

Calling out for help

Help Wanted


Code and Data


Edit Fri Aug 9 09:50:24 2019 anchal:

The last measurement was taken in the wrong way and hence it was mismatched. The ground pin was connected to the wrong position. I retook data today and seems like EOM path and is all healthy :). But the other mismatches still need to be investigated. The Code and Data are updated through git.


Edit Thu Jan 23 19:33:55 2020 anchal:

The mismatched were there because the stage saturates when boost is on. Need to take that measurement with appropriate offset applied through AO Offset. This is done in this CTN:2518.

Attachment 2: NorthFSSInvestigationMismatches.pdf  37 kB  Uploaded Fri Aug 9 17:09:55 2019  | Hide | Hide all
NorthFSSInvestigationMismatches.pdf NorthFSSInvestigationMismatches.pdf NorthFSSInvestigationMismatches.pdf
Attachment 3: NorthFSSInvestigation.pdf  145 kB  Uploaded Fri Aug 9 17:10:07 2019  | Hide | Hide all
NorthFSSInvestigation.pdf NorthFSSInvestigation.pdf NorthFSSInvestigation.pdf NorthFSSInvestigation.pdf NorthFSSInvestigation.pdf NorthFSSInvestigation.pdf NorthFSSInvestigation.pdf NorthFSSInvestigation.pdf NorthFSSInvestigation.pdf NorthFSSInvestigation.pdf NorthFSSInvestigation.pdf NorthFSSInvestigation.pdf NorthFSSInvestigation.pdf
ELOG V3.1.3-